40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 5 17:44:41 2011, kiwamu, Update, General, my plan tonight 
    Reply  Mon Dec 5 19:21:55 2011, kiwamu, Update, LSC, coarse beat note signal : ADC limited above 30 Hz Yarm_ALS_coarse.png
       Reply  Tue Dec 6 00:26:00 2011, kiwamu, Update, LSC, ALS became robust : UGF = 100 Hz 
Message ID: 6074     Entry time: Tue Dec 6 00:26:00 2011     In reply to: 6072
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: ALS became robust : UGF = 100 Hz 

Eventually the instability in the Y end PDH servo turned out to be some kind of an alignment issue.

After carefully realigning the green beam to the Y arm, the UGF of the ALS loop became able to be at more than 50 Hz.

With this UGF it became able to suppress the arm motion to the ADC noise level (few 100 pm in rms).

Now I am scanning the arm length to look for a TEM00 resonance.

 

(the Story)

I have noticed that the spatial fringe pattern of the reflected green light was very sensitive to the pitch motion of ETMY when the green light was locked to the Y arm.

So I realigned the last two launching mirrors to minimize the reflected light. Indeed the misalignment was mainly in the pitch direction.

I basically translated the beam upward by a couple of mm or so.

The amount of the DC reflection is about 2.4 V when it is unlocked and it is now 0.77 mV when the green light is locked.

Quote from #6072

I guess this could be one of the reasons of the unstable behavior in the Y end PDH lock (#6071). (But still it doesn't fully explain the instability).

ELOG V3.1.3-